In a quick return motion mechanism of a Crank and slotted lever type the ratio of maximum velocities is 2.if the length of stroke is 25cm find(1) the length of the slotted lever,



(2)the ratio of time of cutting and written stroke and (3) the maximum cutting velocity per second if the Crank rotates is 30 r.p.n

1
Expert's answer
2022-02-23T13:16:04-0500

"(a)\ \dfrac{x}{25} = 2"


"x = 50\ cm"


"(b)\ \sin\dfrac{25}{50}= \sin({90-\dfrac{\theta}2)}"


"30° = 90 -\dfrac{\theta}2"


"\theta= 120°"


ratio of time of cutting to return strokes = "\dfrac{180-\theta}{\theta}"


"(c)\ 30\ rpm × 2 = 60\ rpm"